OpenStudy (gazela1997):

HELP MEDAL AND FAN 2x^2+x=3-4x

OpenStudy (muzzack):

hiya, r u ready for me to explain now colorfully

OpenStudy (gazela1997):

factor sorry

OpenStudy (gazela1997):

ya sure

OpenStudy (matt101):

First let's collect like terms to simplify things: \[2x^2+x=3-4x\]\[2x^2+5x-3=0\] Then use any method you like to solve for x (e.g. quadratic formula). You should get x = 0.5 and x = -3.

OpenStudy (gazela1997):

but you have to slove this by factoring

OpenStudy (gazela1997):

solve*

OpenStudy (muzzack):

Equation to Solve \(\color{red}{2*x^2+x=3−4*x}\) Solution Steps 1. Negate \(\color{purple}{2*x^2+x=3−4*x}\) 2. Multiply Terms \(\color{blue}{2x^2+x=3−4x}\) 3. Order Terms \(\color{gray}{2x^2+x=-4x+3}\) 4. Move Variables to Left \(\color{lightblue}{2x2+x+4x=3}\) 5. Add Terms \(\color{green}{2x^2+5x=3}\) 6. Factor Trinomial \(\color{darkgreen}{(x+3)*(2x−1)=0}\) Forked Equation 1 \(\color{violet}{x+3=0}\) Forked Equation 2 2x−1=0 Now Solving Equation 1 \(\color{yellow}{x+3=0}\) 7. Move Constants to Right x=0−3 8. Remove Zero Terms x=-3 End of Equation 1 Now Solving Equation 2 2x−1=0 9. Move Constants to Right 2x=0+1 10. Remove Zero Terms 2x=1 11. Divide Both Sides by Term x=12 End of Equation 2 Plug candidate answers into original equation... Answer 1: -3 Answer 2: 12

OpenStudy (matt101):

No problem! I'm going to start with my second equation up there: \[2x^2+5x-3=0\] To factor this, we need to find two numbers that add to give 5 and multiply to give -6 (which you get from 2 * -3). The only two numbers are -1 and 6. We can then rewrite the middle term, 5x, as 6x - x since they're equal: \[2x^2+6x-x-3=0\] Next we need to group the terms in a way that let's us factor them more easily. We can do it like this: \[(2x^2-x)+(6x-3)=0\] I just switched the positions of 6x and -x, and drew brackets around the groups that can now be factored. Now we can finish! \[x(2x-1)+3(2x-1)=0\]\[(2x-1)(x+3)=0\]\[x=\frac{ 1 }{ 2 },-3\]

OpenStudy (muzzack):

yes 1/2 sorry typo